OpenStudy (anonymous):

Write each % as a fraction in Simplest form. 56 1/4%

OpenStudy (lgbasallote):

divide that by 100 then simplify.. for example \[5\text{%} \implies \frac{5}{100}\] then simplify \[\implies \frac{1}{20}\] does that help?

OpenStudy (anonymous):

But what bout the fraction?

OpenStudy (lgbasallote):

that is the fraction..what do you mean?

OpenStudy (anonymous):

56 1/4 &

OpenStudy (anonymous):

1/4!!! Fraction!

OpenStudy (lgbasallote):

oh...

OpenStudy (anonymous):

Say What? I'm in 8th Grade ( 13 yrs) dont understand

OpenStudy (lgbasallote):

\[\Large 4 \frac 12 \text % \implies \frac{4 \frac 12}{100} \implies \frac{\frac 92}{100} \implies \frac{9}{200}\] got it?

OpenStudy (shane_b):

They should put grade levels above the icons :)

OpenStudy (anonymous):

True!

OpenStudy (anonymous):

Help! My math teacher is cuckoo crazy and just gave me this paper with out even explanation! :( I hate Ms. Gonzalez

OpenStudy (lgbasallote):

did you see what i just wrote?

OpenStudy (anonymous):

???? Yes ????? But Idk how to do that?

OpenStudy (lgbasallote):

..i need to go now...i hope @Shane_B can explain further..but just so you know..all i did was improper fractions

OpenStudy (shane_b):

To convert a % to a fraction, you just divide the % by 100. That's what @lgbasallote is saying. You have 56 1/4 %....make that an improper fraction if you need to and then divide that by 100:\[\frac{\frac{225}{4}}{100}=?\]

OpenStudy (lgbasallote):

what i did is very 8th grade level..i assure you